1 and 2 help please. find convergence or non convergence and prove to be true.

2n +1 1. Sn = n 2. Sn = (-1)
Answer #1

ant antat #0 Sn = ant); Smrt = anta+! ht Sn+1-Sn = an+3 – ant - an+34 – lan+l) in to - hint) n -n -1 = 2** +39 - 22-an-nai teInfirmon, le converful to a limit point. & 4 1 them from since А и the sequena has two м (enu! тук.